# Testing strong-field QED with the avalanche precursor

### v1

@@ II.4 NSF OPAL potential
Table 1: Expected parameters of a single Alpha beam at NSF OPAL. The possibility of two such beams at the interaction point is included in the facility design. The value of $a_{0}$ is calculated for the wavelength $\lambda=0.91$ $\mu$m.
@@
Co-pointing precision 1.5 $\mu$m
The anticipated capabilities of NSF OPAL nsf will allow for the exploration of most of the regimes charted in Fig. 1. The relevant parameters are summarized in Table 1. The maximum total power $P_{*}=2P_{0}=50$ PW delivered by two laser pulses, called Alpha beams (each of power $P_{0}$) is marked by an arrow on the top horizontal axis of Fig. 1. Based on the estimates discussed above, the electron-laser interaction can be studied deep in the RD regime, as well as the quantum RR regime. At the peak potential, seed electrons can gain $\chi_{\mathrm{em}}\gtrsim 1$, therefore, the avalanche precursors are within reach.
@@ III Generation of avalanche-type cascades
In this field configuration, we test three realistic options for injecting seed electrons into the focal center: via ionization of a krypton gas, a plasma slab, and a thin aluminium foil. In each case, we identify the development of a cascade with high photon and positron yield. We briefly discuss the temporal evolution of the cascade parameters and compare the target efficiency in terms of the positron yield.
@@ III.1 Targets
Figure 3: Evolution of the cascade parameters for seeding with: (a) a krypton gas, (b) a plasma of density $10n_{c}$, (c) an aluminium foil. For each case, the plots show (from top to bottom): 1) field strength at the focal center [in (a-1), the field strength is shown in gray in arb. units. and we also present ion populations], 2) the number of seed electrons in the simulation box (dashed) and in the focus (solid; the focal volume is defined by $-0.5\lambda\leq x,\,y,\,z\leq 0.5\lambda$); the black markers define the number of seed electrons $N_{\mathrm{seed}}^{\mathrm{foc}}$ in the focus calculated at time $t=-0.5T$, 3) the number of positrons normalized to $N_{\mathrm{seed}}^{\mathrm{foc}}$, 4) $\chi_{e}$ of seed electrons in the focal spot (the light-blue band shows the $90$-percentile of the seed electron distribution in $\chi_{e}$). In (a), we also show the simulation results for an idealized $e^{-}$ target with black solid lines (“$e^{-}$ seed” in the legend), see Section IV.1 for more details. The data is obtained from 3D SMILEI PIC simulations, in which two LP laser pulses counterpropagate, and for each pulse: $a_{0}=550$, focusing $f/2$ (peak power is $P_{0}=25$ PW per pulse), duration 20 fs. The target parameters and results for all cases are summarized in Table 2.
Table 2: Comparison of the parameters for three different targets, and the corresponding positron yield and energy transfer in avalanche-type cascades. The krypton and plasma targets are initially homogeneously distributed in a ball of radius $r=0.75w_{0}\approx 1.5\lambda$. The aluminium foil is shaped as a disk of thickness $d=0.2w_{0}$ along the laser propagation axis, and of radius $r=0.75w_{0}$ in the transverse direction. $N_{\mathrm{seed}}^{\mathrm{foc}}$ is the number of seed electrons in the focus calculated at time $t=-0.5T$. Here, we use $\varepsilon_{\rm laser}\approx 1000$ J to denote the total laser energy, $\varepsilon_{\gamma}$ — the photon energy, and $\varepsilon_{e^{\pm},i}$ — the total energy of electrons, positrons, and ions.
@@
In Fig. 2, we plot the expected ion populations depending on the local field intensity for three noble gases. For 36Kr, the threshold for ionizing the two last $1s$ electrons is below but still close enough to the range of intensities relevant for NSF OPAL. Therefore, in the event of Kr34+ ionization, these electrons will be injected directly into the focal area. For these reasons, we choose krypton as a candidate for seeding avalanches in our setup. We set the gas density to $3.74\times 10^{18}$ cm-3 in the simulation. Assuming a fully ionized gas (initially neutral), this would correspond to the density of $\approx 0.1n_{c}$ for electrons, where $n_{c}=\varepsilon_{0}m\omega^{2}/e^{2}\approx 1.35\times 10^{21}$ cm-3 is the critical plasma density. However, note that this value is not reached at the interaction point, as the electron density rapidly varies in time and space during the interaction, see the details below. As we are interested specifically in the interaction dynamics in the strong-field region, to optimize the simulation time, we use a small ball-shaped target of radius $r=0.75w_{0}\approx 1.34$ $\mu$m. We also consider initially charged ions Kr26+, as outer electrons are extracted at earlier stages and do not participate in the interaction in the focal center. [^5] We performed additional simulations to confirm this. Simulations with an initially neutral target and with a pre-charged target result in the same number of seed electrons occurring in the focal center of the laser field. Let us mention that our simulations show that, potentially, argon can also be used for triggering cascades. However, when compared to krypton, the number of seed electrons injected in the focal area will be lower for the studied here setup parameters. Using xenon could be efficient at higher intensities.
Next, we consider plasma seeding, and take as well a small ball-shaped target of radius $r=0.75w_{0}\approx 1.34$ $\mu$m, and set the plasma density to $10n_{c}$. [^6] The possibility of pre-generating such plasma at the interaction point is a nontrivial question, however, out of the scope of this work. Note that under the considered parameters, the total laser field remains unaffected during the interaction with such a plasma target.
Finally, for the aluminium foil, we assume that the target is a thin disk of thickness $d=0.2w_{0}\approx 0.356$ $\mu$m and $r=0.75w_{0}\approx 1.34$ $\mu$m in the transverse direction. The laser pulses are normally incident on the foil surface. The targets’ parameters are summarized in Tab. 2. Note that the choice of the target thickness satisfies the condition $a_{0}>\pi(n_{0}/n_{c})(d/2)/\lambda\approx 350$, ensuring that two pulses can penetrate the target from its sides. [^7] This condition on the target thickness and density implies that a single laser pulse of strength $a_{0}$ can separate (i.e. push) all of the electrons from ions Vshivkov et al. (1998); Bulanov et al. (2016).
@@ III.2 Simulation setup
Let us briefly describe our numerical setup. Each laser pulse is strongly focused at $\mathbf{r}=\mathbf{0}$ and propagates along either the positive or negative $x$-direction. The field wavelength is set to $\lambda=0.91$ $\mu$m throughout and the corresponding period is indicated as $T=\lambda/c$. For each pulse, the temporal envelope is Gaussian with duration $\tau_{\mathrm{FWHM}}=20$ fs (per intensity profile), so that the components of the field are $\propto a_{0}\exp(-t^{2}/\tau^{2})$, where $\tau=\tau_{\mathrm{FWHM}}/\sqrt{2\ln 2}$. The spatial profile in the focal plane is also Gaussian in the transverse direction with waist $w_{\mathrm{FWHM}}$ (intensity profile), so that the field components are $\propto a_{0}\exp(-r^{2}/w_{0}^{2})$, where $w_{0}=w_{\mathrm{FWHM}}/\sqrt{2\ln 2}$. In this Section, we consider $f/2$ focusing, so $w_{\mathrm{FWHM}}=2.3$ $\mu$m ($w_{0}\approx 1.96\lambda$). The power is fixed to $P_{0}=25$ PW per pulse and the polarization is linear along the $y$-direction, so the maximum field amplitude is $a_{0}=550$ and the corresponding intensity $I_{0}\approx 5\times 10^{23}$ W/cm2 (hence, for the total field, $P_{*}=50$ PW, $a_{*}=1100$, and intensity $I_{\mathrm{total}}\approx 2\times 10^{24}$ W/cm2). Under these parameters, a single laser pulse has an energy of $\approx 500$ J.
@@
We select the following settings for our 3D simulations with the SMILEI PIC code.Derouillat et al. (2018) The simulation box is sized $20\lambda$ in the $x$-direction (along the laser pulse propagation) and $40\lambda$ in the $y$- and $z$-direction. The laser field focus and the target are located at the box center at the origin $\mathbf{r}=\mathbf{0}$. The time $t=0$ is chosen as the moment when the laser envelope centers reach $\mathbf{r}=\mathbf{0}$. At the initial stage of each simulation, the target is not exposed to the laser field. The pulses are injected from the box boundaries and are incident on the target from the sides during the run. As we define the field profile in the focal plane, to ensure consistency with Maxwell’s equations, we use a numerical procedure for backward propagation of the laser field to the boundaries before each simulation. For more details, see Ref.Thiele, Skupin, and Nuter (2016) and the LaserOffset option description in the SMILEI code manual.smi
@@ III.3 Injection of electrons and the cascade onset
A cascade can be triggered only by seed electrons located near the focal center. Let us define the focal volume by a box of volume $\lambda^{3}$ at the center and track the number of electrons in it. This number is shown by the red solid line in Fig. 3(a-2). After a sharp increase, the number of seed electrons in the focus gradually decreases as they are pushed by the field. The two bumps near $t\approx-4.5T$ and $-0.5T$ correspond to the ionization of the $1s$ electrons in the focus. The maximum and mean $\chi_{e}$ for electrons in the focus exceed unity [see Fig. 3(a-4)]. Therefore, such electrons can trigger an avalanche-type cascade.
@@
It is convenient to introduce the number of seed electrons in the focus $N_{\mathrm{seed}}^{\mathrm{foc}}$ calculated at a fixed time close to the cascade onset. This allows for a consistent measure of different cascade parameters per seed electrons. We choose $t=-0.5T$, as at this moment the laser field amplitude is close to its maximum, and the conditions are favourable for the cascade development. We explicitly show $N_{\mathrm{seed}}^{\mathrm{foc}}$ for each target in Figs. 3(a-2), (b-2), and (c-2) and present the corresponding values in Tab. 2. Because $N_{\mathrm{seed}}^{\mathrm{foc}}$ is proportional to the initial target density, a foil provides the highest value.
@@ III.4 Positron yield and energy transfer
Figure 4: Time dependence of the energy balance between the laser field, radiation, and particles. The results are obtained from PIC simulations with an aluminium foil target (see also Tab. 2). The data is normalized to the total energy of two laser pulses ($\approx$1000 J). The orange curve is obtained by integrating the spectral data for photons with energy $>1$ MeV. The laser field configuration is described in Section III.2.
@@ III.5 Final particle distributions
Figure 5: The final particle spectra. In all plots, the black line shows the spectrum calculated for all particles in the simulation box, the orange line corresponds to particles propagating in the polar angle opening $0.35\pi<\theta<0.65\pi$, the blue line to particles propagating in $0<\theta<0.15\pi$ (recall that $\theta=0$ corresponds to the laser propagation axis). The green dashed line in (c) shows the spectrum exclusively for electrons that were created in the cascade. The vertical dark-blue and magenta lines show the maximum $e^{-}$ energy estimated from Eq. (2) and $\gamma_{\mathrm{em}}mc^{2}$ given by Eq. (4), respectively (see also Fig. 1). All spectra are normalized to the total number of the corresponding species. The injection is implemented via a krypton gas (see Tab. 2), and the laser field configuration and parameters are presented in Section III.2.
Figure 6: (a) Final angular distribution of photons (in arb. units). Note that the electric field is polarized linearly in the $y$-direction ($\theta=\pi/2$, $\varphi=0$ or $\pi$). (b-d) Final photon and positron angular-energy distributions (normalized to unity). The parametrisation of angles is illustrated in (a), which is used throughout. The green curves show the energy-integrated angular distribution on a linear scale in arb. units. (namely, the 2D plot integrated along the $y$-axis). Here, the target is krypton gas (see Tab. 2), and the laser field configuration and parameters are presented in Section III.2.
The final energy and angular distributions resulting from the interaction are fundamental quantities of interest for experiments. The normalized final particle distributions resulting from the three considered targets share similar features. Therefore, in this subsection, we provide the simulation data obtained for the krypton gas. The final particle spectra and angular-energy distributions are presented in Figs. 5 and 6.
@@
We now consider the angular dependence of the spectra from this perspective. Figure 6 shows the final angular-resolved distributions. The choice of the azimuthal $-\pi<\varphi\leq\pi$ and polar $0\leq\theta<\pi$ angles is illustrated in Fig. 6(a). Recall that the laser pulses propagate along the positive or negative $x$-direction ($\theta=0$ or $\theta=\pi$, respectively) and the electric field is polarized in the $y$-axis ($\theta=\pi/2$, $\varphi=0$ or $\pi$). For brevity, we refer to these directions as longitudinal and transverse, respectively.
As seen from Figs. 6(a), (b), and (d), photons are predominantly emitted transversely in the $xy$-plane, and the most energetic photons are distributed around the transverse direction. For illustration, in Fig. 5(a), we also show the spectra for photons emitted around a selected direction with the angular spread $\Delta\theta=0.15\pi$: longitudinally, so that $0<\theta<\Delta\theta$ (the blue curve), and transversely, $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$ (the orange curve).
@@
Let us consider the high-energy tails for the transversely propagating particles in more detail [see the orange curves in Figs. 5(b) and (c)]. We can compare the energy of the tails to the estimates based on the RR models introduced in Sections II.1 and II.2. The classical expression with the LL term, given by Eq. (2), [^10] As we mentioned, Eq. (2) is derived for a rotating electric field, but it can be used in estimates for other field configurations. predicts a maximum electron energy $\varepsilon_{e}=mc^{2}\gamma_{e}\approx 270$ MeV for $a_{*}=1100$ and $\lambda=0.91$ $\mu$m. We mark this threshold by a blue vertical line in Figs. 5(b) and (c). As one may see, the exponential tails extend beyond the prediction of the classical theory, hence, as expected, the RR is manifestly quantum. The exponential cut-off location can be estimated with Eq. (4), which gives $\gamma_{\mathrm{em}}mc^{2}\approx 360$ MeV (also shown in the figure with the magenta line). This is coherent with our simulation results.
Summarizing, most of the radiation and new particles generated in the cascade are ejected in the direction of the electric field polarization axis. The spectra measured at the corresponding angles can used to characterize the interaction in the focal center.

### v2

@@ II.4 NSF OPAL potential
[…]
@@
[…]
The anticipated capabilities of NSF OPAL nsf (b) will allow for the exploration of most of the regimes charted in Fig. 1. The relevant parameters are summarized in Table 1. The maximum total power $P_{*}=2P_{0}=50$ PW delivered by two laser pulses, called Alpha beams (each of power $P_{0}$) is marked by an arrow on the top horizontal axis of Fig. 1. Based on the estimates discussed above, the electron-laser interaction can be studied deep in the RD regime, as well as the quantum RR regime. At the peak potential, seed electrons can gain $\chi_{\mathrm{em}}\gtrsim 1$, therefore, the avalanche precursors are within reach.
@@ III Generation of avalanche-type cascades
Figure 2: Scheme of the interaction: two incident LP laser pulses (with the parameters listed in Tab. 1) form a standing wave and ionize a krypton gas. Electrons are injected into the laser focus and trigger an avalanche-type cascade. The generated $e^{\pm}$ and $\gamma$-photons outgoing from the interaction point are depicted with blue and yellow spheres, respectively. $\Delta\theta$ denotes a polar angle opening (see Sec. III.5).
@@
In this field configuration, we test three realistic options for injecting seed electrons into the focal center: via ionization of a krypton gas, a plasma slab, and a thin aluminium foil. The interaction scheme is illustrated in Fig. 2 for the former case. For each target, we identify the development of a cascade with high photon and positron yield. We briefly discuss the temporal evolution of the cascade parameters and compare the target efficiency in terms of the positron yield.
@@ III.1 Targets
Table 2: Comparison of the parameters for three different targets, and the corresponding positron yield and energy transfer in avalanche-type cascades. The krypton and plasma targets are initially homogeneously distributed in a ball of radius $r=0.75w_{0}\approx 1.5\lambda$. The aluminium foil is shaped as a disk of thickness $d=0.2w_{0}$ along the laser propagation axis, and of radius $r=0.75w_{0}$ in the transverse direction. $N_{\mathrm{seed}}^{\mathrm{foc}}$ is the number of seed electrons in the focus calculated at time t=0.5Tt=-0.5T. Here, we use $\varepsilon_{\rm laser}\approx 1000$ J to denote the total laser energy, $\varepsilon_{\gamma}$ — the photon energy, and $\varepsilon_{e^{\pm},i}$ — the total energy of electrons, positrons, and ions.
@@
In Fig. 3, we plot the expected ion populations depending on the local field intensity for three noble gases. To obtain these curves, we used the Perelomov-Popov-Terent’ev formula for strong-field ionization derived for small values of the Keldysh parameter Perelomov, Popov, and Terent’ev (1966, 1967); Perelomov and Popov (1967) (see also Ref.Ammosov, Delone, and Krainov (1986)). In the version of SMILEI PIC code used here, this formula is implemented following the algorithm suggested in Ref.Nuter et al. (2011) Let us note that we do not take into account barrier suppression corrections, which are expected to be important for the dynamics of outer shell ionization by an incident laser pulse.Artemenko and Kostyukov (2017); Ouatu et al. (2022); Mironov et al. (2025) As we suggest below, in the parameter range of NSF OPAL, it is beneficial to use the $1s$-subshell of krypton as a source of seed electrons triggering a cascade. Ionization of these atomic states appears insensitive to barrier suppression or other corrections to the strong-field ionization model.Mironov et al. (2025)
[…]
To conclude the discussion about targets, let us estimate the relevance of QED showers linking bremsstrahlung and Bethe-Heitler processes in the targets. For this, we use a simple estimate of the radiation lengthBethe and Heitler (1934) $X_{0}=1/[4\alpha r_{e}^{2}n_{0}Z(Z+1)\log(183Z^{-1/3})]$ (see also Ref.Pouyez et al. (2025b) for recent progress on the theory of showers in matter). Under normal conditions, one obtains for aluminium ($Z=13$, $n_{0}\approx 1.6\times 10^{23}$ cm-3) $X_{0}\sim 3$ cm and orders magnitude larger for krypton. Hence, by using smaller-sized targets, it is possible to exclude the development of showers generated by the bremsstrahlung and Bethe-Heitler processes, which is achievable in future experiments.
@@ III.2 Simulation setup
Figure 4: Distribution of the $E_{y}$ component of the field of two counterpropagating laser pulses in a plane $z=0$ before the interaction. The green circle shows the initial shape of the krypton and $10n_{c}$-plasma targets used in simulation (see also Tab. 2).
Figure 5: Evolution of the cascade parameters for seeding with: (a) a krypton gas, (b) a plasma of density $10n_{c}$, (c) an aluminium foil. For each case, the plots show (from top to bottom): 1) field strength at the focal center [in (a-1), the field strength is shown in gray in arb. units. and we also present ion populations], 2) the number of seed electrons in the simulation box (dashed) and in the focus (solid; the focal volume is defined by $0.5\lambda\leq x,\,y,\,z\leq 0.5\lambda$); the black markers define the number of seed electrons $N_{\mathrm{seed}}^{\mathrm{foc}}$ in the focus calculated at time t=0.5Tt=-0.5T, 3) the number of positrons normalized to $N_{\mathrm{seed}}^{\mathrm{foc}}$, 4) $\chi_{e}$ of seed electrons in the focal spot (the light-blue band shows the $90$-percentile of the seed electron distribution in $\chi_{e}$). In (a), we also show the simulation results for an idealized $e^{-}$ target with black solid lines ($e^{-}$ seed” in the legend), see Section IV.1 for more details. The data is obtained from 3D SMILEI PIC simulations, in which two LP laser pulses counterpropagate, and for each pulse: $a_{0}=550$, focusing $f/2$ (peak power is $P_{0}=25$ PW per pulse), duration 20 fs. The target parameters and results for all cases are summarized in Table 2.
[…]
The field components are calculated by solving Maxwell’s equations numerically with the method suggested in Ref.Thiele, Skupin, and Nuter (2016) We specify the transverse magnetic field components in the focal plane, so that $B_{y}|_{x=0}=0$ and the $B_{z}$ component for the pulse propagating in the positive $x$ direction reads (in the dimensionless units):
$B_{z}|_{x=0}=a_{0}e^{-(y^{2}+z^{2})/w_{0}^{2}}e^{-(t/\tau)^{2}}\cos t,$ (7)
We flip the sign of $B_{z}|_{x=0}$ for the counterpropagating pulse, so that $x=0$ corresponds to the electric antinode center of the total standing-wave field. Before running a full simulation of the interaction with a target, the field of each laser pulse is propagated backwards from the focal plane to the corresponding simulation box boundary. This provides boundary conditions ensuring the field shape at the focus. Then, all the field components are evaluated in the PIC loop at all time-space points during the simulation run. See also the LaserOffset option description in the SMILEI code manual.smi A snapshot of the numerically calculated $E_{y}$ component distribution before the interaction is shown in Fig. 4.
The expected contrast at NSF OPAL is at the level of $10^{10}$ in the far field up to 100 ps before the main pulse (the facility design is based on optical parametric chirped-pulse amplification technique). We assume that under these parameters, the target (in particular, a foil) survives the interaction with the prepulse, namely, it does not expand significantly before the arrival of the main pulses. Therefore, in this work, we do not simulate the interaction with the pedestal and consider only the main femtosecond-scale pulses. Notably, in the double-pulse geometry, the electrons are compressed by the ponderomotive pressure from the two sides of the target, which also helps to keep the target confined before the main pulse arrival.
@@
We set PIC simulation parameters as follows. The simulation box is sized $20\lambda$ in the $x$-direction (along the laser pulse propagation) and $40\lambda$ in the $y$- and $z$-direction. The laser field focus and the target are located at the box center at the origin $\mathbf{r}=\mathbf{0}$. The time $t=0$ is chosen as the moment when the laser envelope centers reach $\mathbf{r}=\mathbf{0}$. At the initial stage of each simulation, the target is not exposed to the laser field. The laser pulses are injected from the box boundaries and are incident on the target from the sides during the run as described above.
@@ III.3 Injection of electrons and the cascade onset
A cascade can be triggered only by seed electrons located near the focal center. Let us define the focal volume by a box of volume $\lambda^{3}$ at the center and track the number of electrons in it. This number is shown by the red solid line in Fig. 5(a-2). After a sharp increase, the number of seed electrons in the focus gradually decreases as they are pushed by the field. The two bumps near $t\approx-4.5T$ and 0.5T-0.5T correspond to the ionization of the $1s$ electrons in the focus. The maximum and mean $\chi_{e}$ for electrons in the focus exceed unity [see Fig. 5(a-4)]. Therefore, such electrons can trigger an avalanche-type cascade.
@@
It is convenient to introduce the number of seed electrons in the focus $N_{\mathrm{seed}}^{\mathrm{foc}}$ calculated at a fixed time close to the cascade onset. This allows for a consistent measure of different cascade parameters per seed electrons. We choose t=0.5Tt=-0.5T, as at this moment the laser field amplitude is close to its maximum, and the conditions are favourable for the cascade development. We explicitly show $N_{\mathrm{seed}}^{\mathrm{foc}}$ for each target in Figs. 5(a-2), (b-2), and (c-2) and present the corresponding values in Tab. 2. Because $N_{\mathrm{seed}}^{\mathrm{foc}}$ is proportional to the initial target density, a foil provides the highest value.
@@ III.4 Positron yield and energy transfer
Figure 6: Time dependence of the energy balance between the laser field, radiation, and particles. The results are obtained from PIC simulations with an aluminium foil target (see also Tab. 2). The data is normalized to the total energy of two laser pulses ($\approx$1000 J). The orange curve is obtained by integrating the spectral data for photons with energy >1>1 MeV. The laser field configuration is described in Section III.2. Note that the figure shows the total energy contained in the actual simulation box. The maximum laser energy (as well as the energy balance) is reached at $t\gtrsim-5T$, when the pulses fully enter the simulation box.
@@ III.5 Final particle distributions
Figure 7: The final spectra of (a) photons, (b) electrons, and (c) positrons. In all plots, the black line shows the spectrum calculated for all particles in the simulation box, the orange line corresponds to particles propagating in the polar angle opening $0.35\pi<\theta<0.65\pi$, the blue line to particles propagating in $0<\theta<0.15\pi$ (recall that $\theta=0$ corresponds to the laser propagation axis). The green dashed line in (c) shows the spectrum exclusively for electrons that were created in the cascade. The vertical dark-blue and magenta lines show the maximum $e^{-}$ energy estimated from Eq. (2) and $\gamma_{\mathrm{em}}mc^{2}$ given by Eq. (4), respectively (see also Fig. 1). All spectra are normalized to the total number of the corresponding species. The injection is implemented via a krypton gas (see Tab. 2), and the laser field configuration and parameters are presented in Section III.2.
Figure 8: (a) Final angular distribution of photons (in arb. units). Note that the electric field is polarized linearly in the $y$-direction ($\theta=\pi/2$, $\varphi=0$ or $\pi$). (b-e) Final photon and positron angular-energy distributions (normalized to unity). The parametrisation of angles is illustrated in (a), which is used throughout. The green curves show the energy-integrated angular distribution on a linear scale in arb. units. (namely, the 2D plot integrated along the vertical axis). Here, the target is krypton gas (see Tab. 2), and the laser field configuration and parameters are presented in Section III.2.
The final energy and angular distributions resulting from the interaction are fundamental quantities of interest for experiments. The normalized final particle distributions resulting from the three considered targets share similar features. Therefore, in this subsection, we provide the simulation data obtained for the krypton gas. The final particle spectra and angular-energy distributions are presented in Figs. 7 and 8.
@@
We now consider the angular dependence of the spectra from this perspective. Figure 8 shows the final angular-resolved distributions. The angles are chosen with the following convention (illustrated in Fig. 8a): $\varphi$ is the azimuthal angle, $\pi<\varphi\leq\pi$, counted from the positive direction of the $y$-axis in the $(yz)$-plane, and $\theta$ is the polar angle, $0\leq\theta<\pi$, counted from the positive direction of the $x$-axis. Recall that the laser pulses propagate along the positive or negative $x$-direction ($\theta=0$ or $\theta=\pi$, respectively) and the electric field is polarized in the $y$-axis ($\theta=\pi/2$, $\varphi=0$ or $\pi$). For brevity, we refer to these directions as longitudinal and transverse, respectively.
As seen from Figs. 8(a), (b), and (d), photons are predominantly emitted transversely in the $xy$-plane, and the most energetic photons are distributed around the transverse direction. For illustration, in Fig. 7(a), we also show the spectra for photons emitted around a selected direction with the angular spread $\Delta\theta=0.15\pi$: longitudinally, so that $0<\theta<\Delta\theta$ (the blue curve), and transversely, $\pi/2-\Delta\theta<\theta<\pi/2+\Delta\theta$ (the orange curve; see also Fig. 2 for illustration). Note that the value $\Delta\theta=0.15\pi$ is chosen for illustration, and variation of $\Delta\theta$ does not affect the overall discussion presented below as long as the transverse and longitudinal motion can be separated (see Appendix C and Fig. 12 therein for more details).
@@
[…]
Summarizing, most of the radiation and new particles generated in the cascade are ejected in the direction of the electric field polarization axis. The spectra measured at the corresponding angles can be used to characterize the interaction in the focal center.